OpenStudy (itsbribro):

what is 4x – 3y = 24 in a graph?

OpenStudy (whpalmer4):

Get out your graph paper and a ruler. And a pencil :-) Set x=0, and solve for y. Make a mark at (0,y). Set y=0, and solve for x. Make a mark at (x,0). Now draw a straight line through those two points. Congratulations, you have a graph of \(4x-3y=24\) suitable for framing :-)

OpenStudy (whpalmer4):

Yes, your graph will look like one of those.

OpenStudy (itsbribro):

im asking which one i dont know and my couisn needs it

OpenStudy (whpalmer4):

Do as I suggested, but skip making the graph. Just find (0,y) and (x,0). Look at each answer choice and find the one that goes through those two points.

OpenStudy (whpalmer4):

\[4x-3y=24\]set \(x=0\) \[4(0)-3y=24\]\[0-3y=24\]\[-3y=24\]\[y=\]

OpenStudy (whpalmer4):

for the other point: \[4x-3y=24\] set \(y=0\) \[4x-3(0) = 24\]\[4x-0=24\]\[4x=24\]\[x=\]